# Near-degeneracy effects in Quadrupolar Mixed Modes

### v1

@@ 1 Introduction
Understanding the internal dynamics of evolved solar-like stars remains one of the major challenges in asteroseismology. Internal dynamical processes can be probed using oscillations that consist of acoustic pressure modes (p modes) and buoyancy-driven gravity modes (g modes). In evolved solar-like stars, p and g modes can couple across their respective cavities, giving rise to mixed modes (e.g. 1989nos..book.....U; 2016PASJ...68...91T), first detected by 2011Sci...332..205B. Mixed-mode analyses have revealed the core and the envelope rotation rates of subgiant and red giant stars (e.g. 2011Sci...332..205B; 2012ApJ...756...19D; 2012A&A...548A..10M; 2014A&A...564A..27D; 2016ApJ...817...65D; 2018A&A...616A..24G; 2024A&A...688A.184L). However, measured rotational gradients are substantially weaker than those predicted by hydrodynamical theoretical models of angular momentum evolution (e.g. 2012A&A...544L...4E; 2013A&A...549A..74M; 2013A&A... [^5] A..54C; 2017A&A...599A..18E; 2019A&A...631L...6E). This discrepancy indicates the presence of missing, angular momentum transport mechanisms in our representation of angular momentum transport in stars. Several processes have been proposed to enhance transport, including the meridional circulation (1992A&A...265..115Z; 1998A&A...334.1000M; 2004A&A...425..229M; 2009A&A...495..271D), hydrodynamical instabilities (1992A&A...265..115Z), mixed modes (2015A&A...579A..30B; 2025A&A...699A.310B), internal gravity waves (1993A&A...279..431S; 2014ApJ...796...17F; 2017A&A...605A..31P; 2025ApJ...983L..38R), and magnetic processes such as transport by Alfvén waves and magnetic tension (e.g. 1987MNRAS.226..123M; 2007A&A...474..145Z; 2022A&A...661A.119G; 2021A&A...646A..19T), magneto-rotational instabilities (2024A&A...683A..12M), magnetic webs (2025ApJ...989L...4S), or transport by the Tayler-Spruit dynamo (e.g. 1973MNRAS.161..365T; 2002A&A...381..923S; 2019MNRAS.485.3661F; 2023A&A...677A...6M; 2025ApJ...988..195S). To date, no observational constraints have unequivocally identified any overarching dominant physical mechanism responsible for angular momentum transport along stellar evolution.
@@ 2.1.1 The general rotating eigenvalue problem
where $\hat{\mathcal{L}}$ denotes the linear operator that encapsulates the perturbations of the equilibrium due to volumetric forces and acts on the displacement field $\@vec{\delta r}$ associated with stellar oscillations. The solutions of this eigenvalue problem are given by a set of eigenvalues $\omega_{0,n,\ell,m}$ and eigenfunctions $\@vec{\delta r}_{0,n,\ell,m}$ that can be decomposed on the spherical harmonics basis as:
$\begin{split}\@vec{\delta r}_{0,n,\ell,m}&=\xi_{0,r,\ell,n,m}(r)Y_{\ell}^{m}(\theta,\phi)\@vec{e}_{r}\\ &+\xi_{0,h,\ell,n,m}(r)\left(\frac{\partial Y_{\ell}^{m}}{\partial\theta}\@vec{e_{\theta}}+\frac{1}{\sin(\theta)}\frac{\partial Y_{\ell}^{m}}{\partial\varphi}\@vec{e_{\varphi}}\right)\,.\end{split}$ (2)
@@
$\begin{split}\@vec{\delta r}_{0,i}&=\xi_{0,r,i}(r)Y_{\ell}^{m}(\theta,\phi)\@vec{e}_{r}\\ &+\xi_{0,h,i}(r)\left(\frac{\partial Y_{\ell}^{m}}{\partial\theta}\@vec{e_{\theta}}+\frac{1}{\sin(\theta)}\frac{\partial Y_{\ell}^{m}}{\partial\varphi}\@vec{e_{\varphi}}\right)\,.\end{split}$ (3)
@@
When including stellar rotation into the movement equation, the spherical-symmetry-induced degeneracy is lifted and the solutions to the eigenvalue problem depend on $m$. At leading order in $\Omega$, the angular rotation rate of the star, the eigenvalue problem becomes (e.g. 2010aste.book.....A; 2017A&A...605A..75D; 2025arXiv250926319A):
@@
$\omega^{2}\mathbb{I}\@vec{a}-\omega R\@vec{a}-D\@vec{a}=0\,$, (8)
where $\mathbb{I}$ is the identity matrix, $D=\mathrm{diag}(\omega_{0,i}^{2})$, a\@vec{a} is the vector of the $a_{i}$ (i.e. the coefficients of the decomposition of $\@vec{\delta r}$ in the unperturbed base following Eq. (5)) and:
@@ 2.2 First order symmetric rotational effects: diagonal terms
may be asymptotically expressed using the seismic $\zeta$ function. For a given unperturbed mixed mode, the asymptotic expression of the $\zeta$-function provides the ratio between the mode inertia in the g-mode cavity and the total mode inertia (2013A&A...549A..75G):
@@ 2.3 Near-degeneracy effects: non-diagonal terms
$\gamma_{c,ij}=\int_{\mathrm{g}}K_{i,j}(r)\mathrm{d}r\,$, (17)
@@
$\gamma_{e,ij}=\int_{\mathrm{p}}K_{i,j}(r)\mathrm{d}r\;.$ (18)
@@ 2.3.1 New asymptotic formula for the near degeneracy effects
Using the change of variable111it is possible because in the g-mode cavity $\frac{N}{r}>0$ hence the integral is a strictly increasing function of $r$
@@ 3.1 Description of the Bayesian fitting procedure
To accomplish this, we employ a Bayesian framework to fit the oscillation modes in the PSD. The p- and mixed-mode patterns are created using the sloscillations code developed in 2019MNRAS.488..572K, extended to incorporate the near-degeneracy effects based on Section 2. The various asymptotic formulations used to produce the mode pattern are presented below.
@@ 3.1.1 Pure pressure and gravity modes
$\nu_{{\rm g}_{n,\ell}}=\frac{1}{\Delta\Pi_{\ell}(n+\frac{\ell}{2}+\varepsilon_{\rm g})}\,$, (33)
@@ 3.1.2 Mixed modes with rotation
Mixed mode frequencies are computed using equation (1979PASJ...31...87S):
@@
Then we introduce the effects of rotation at first perturbative order for the $\ell=1$ components and solve the full rotating quadratic eigenvalue problem with asymptotic matrix elements for $\ell=2$ (based on our asymptotic formulation Eq. (26)) modes using the slepc4py (DALCIN20111124; 10.1145/1089014.1089019) Python library.
@@ 3.1.3 Modelling the linewidth, amplitudes and heights of the modes
This paragraph describes the equations and methodologies used to compute the linewidths amplitudes and heights of the modes in the sloscillations code by 2019MNRAS.488..572K for the sake of completeness and clarity. We adopt effective-temperature-dependent scaling relations for the p mode linewidths, parameterised as a function of temperature, following the empirical prescriptions of corsaro_asteroseismology_2012; 2017ApJ...835..172L. Radial-mode amplitudes follow a Gaussian envelope centred on $\nu_{\rm max}$ with width tied to the envelope full width at half maximum and a peak amplitude that scales with global parameters ($\nu_{\rm max}$, $\Delta\nu$). The per-degree amplitudes are set by standard photometric visibilities as in 2019MNRAS.488..572K and, for non-radial multiplets, by geometric inclination ($\theta_{\mathrm{inc}}$) factors (2003ApJ...589.1009G). For mixed $\ell=1$ and $\ell=2$ modes, we modulate the nominal p mode amplitudes by $(1-\zeta)^{\frac{1}{2}}$ (2018A&A...618A.109M; Basu_Chaplin_2017). Mode heights in the power spectrum are computed from the time-domain amplitudes and the linewidths using the finite-observation-length correction of Fletcher_2006. For mixed modes, we use the mixed-mode generalisation that explicitly includes the inertia ratio $\zeta$ (Basu_Chaplin_2017), ensuring that g-dominated mixed modes appear with appropriately reduced height at fixed amplitude. Each mode contributes a Lorentzian profile to the PSD, centred at the asymptotic frequency and with the height and width described above.
The background is modelled with the usual super-Lorentzian granulation components and white noise (e.g 2014A&A...570A..41K). It is generated in the time domain via a Gaussian-process (GP) representation of convection and granulation following 2019MNRAS.488..572K, where a kernel encapsulates the covariance of stochastic surface motions (e.g. 2017AJ....154..220F; 2019MNRAS.489.5764P; O’Sullivan_Aigrain_2024). This data is Fourier transformed, and the resulting PSD provides the granulation floor plus white noise that forms the background in the PSD. On top of this background, the oscillation spectrum is added, producing a self-consistent synthetic PSD where both the convective noise and p-mode power originate from the same underlying time-domain process.
@@ 3.1.4 Priors on p modes and $\ell=1$ mixed modes
3.2 Priors on the $\ell=2$ period spacing $\Delta\Pi_{2}$ and phase $\varepsilon_{g,2}$
@@
3.3 Priors on the $\ell=2$ modes coupling parameter $q_{2}$
We examine what constraints can be put on $q_{2}$, the coupling factor of $\ell=2$ modes, depending on the inclination angle of the star compared to the line of sight ($\theta_{\mathrm{inc}}$) (which we estimate by the preliminary fit of the $\ell=1$ modes – Appendix C). When the $\ell=2,\,m=0$ modes are visible, we use the technique described in Sec. 3.3.1. When the angle of inclination makes the $\ell=2$ and $m=0$ modes not visible, we use the result of Sec. 3.3.2.
3.3.1 Using the $m=0$ modes
@@
3.3.2 Using the $\ell=1$ mode coupling parameter $q_{1}$
@@
3.4 Adopted fitting strategy
@@ 4.2 Results of the Bayesian asymptotic fitting
$\varepsilon$ $0.257\pm 0.004$ $0.422\pm 0.007$
$\delta\nu_{0,1}$ (µHz) $0.118\pm 0.02$ $0.345\pm 0.014$
$\delta\nu_{0,2}$ (µHz) $3.411\pm 0.021$ $2.187\pm 0.025$
@@
$\Delta\Pi_{1}$ (s) $111.181\pm 0.025$ $97.010\pm 0.007$
$q_{1}$ $0.378\pm 0.001$ $0.194\pm 0.001$
$\varepsilon_{g,1}$ $0.310\pm 0.006$ $0.280\pm 0.003$
$\Delta\Pi_{2}$ (s) $64.421\pm 0.014$ $56.022\pm 0.014$
$q_{2}$ $040\pm 0.001 0.0$$019\pm 0.001$
$\varepsilon_{g,2}$ $0.138\pm 0.008$ $0.129\pm 0.015$
${\Omega_{\mathrm{core}}}/{2\pi}$ (nHz) $781\pm 4$ $677\pm 4$
${\Omega_{\mathrm{env}}}/{2\pi}$ (nHz) $53\pm 6$ $43\pm 6$
$\theta_{\mathrm{inc}}$ (°) $85$ (fixed) $52\pm 3$

### v2

@@ 1 Introduction
Understanding the internal dynamics of evolved solar-like stars remains one of the major challenges in asteroseismology. Internal dynamical processes can be probed using oscillations that consist of acoustic pressure modes (p modes) and buoyancy-driven gravity modes (g modes). In evolved solar-like stars, p and g modes can couple across their respective cavities, giving rise to mixed modes (e.g. 1989nos..book.....U; 2016PASJ...68...91T), first detected by 2011Sci...332..205B. Mixed-mode analyses have revealed the core and the envelope rotation rates of subgiant and red giant stars (e.g. 2011Sci...332..205B; 2012ApJ...756...19D; 2012A&A...548A..10M; 2014A&A...564A..27D; 2016ApJ...817...65D; 2018A&A...616A..24G; 2024A&A...688A.184L). However, measured rotational gradients are substantially weaker than those predicted by hydrodynamical theoretical models of angular momentum evolution (e.g. 2012A&A...544L...4E; 2013A&A...549A..74M; 2013A&A... [^5] A..54C; 2017A&A...599A..18E; 2019A&A...631L...6E). This discrepancy indicates the presence of missing, angular momentum transport mechanisms in our representation of angular momentum transport in stars. Several processes have been proposed to enhance transport, including the meridional circulation (1992A&A...265..115Z; 1998A&A...334.1000M; 2004A&A...425..229M; 2009A&A...495..271D), hydrodynamical instabilities (1992A&A...265..115Z), mixed modes (2015A&A...579A..30B; 2025A&A...699A.310B), internal gravity waves (1993A&A...279..431S; 2014ApJ...796...17F; 2017A&A...605A..31P; 2025ApJ...983L..38R), and magnetic processes such as transport by Alfvén waves and magnetic tension (e.g. 1987MNRAS.226..123M; 2007A&A...474..145Z; 2021A&A...646A..19T; 2022A&A...661A.119G), magneto-rotational instabilities (2024A&A...683A..12M), magnetic webs (2025ApJ...989L...4S), or transport by the Tayler-Spruit dynamo (e.g. 1973MNRAS.161..365T; 2002A&A...381..923S; 2019MNRAS.485.3661F; 2023A&A...677A...6M; 2025ApJ...988..195S). To date, no observational constraints have unequivocally identified any overarching dominant physical mechanism responsible for angular momentum transport along stellar evolution.
@@ 2.1.1 The general rotating eigenvalue problem
where $\hat{\mathcal{L}}$ denotes the linear operator that encapsulates the perturbations of the equilibrium due to volumetric forces and acts on the displacement field $\@vec{\delta r}$ associated with stellar oscillations:
$\hat{\mathcal{L}}\,\@vec{\delta r}=\frac{\nabla p^{\prime}}{\rho}+\nabla\Phi^{\prime}-\frac{\rho^{\prime}}{\rho^{2}}\,\nabla p\,.$ (2)
where $\rho$ and $p$ are the equilibrium pressure and density of the star and $p^{\prime},\rho^{\prime},\Phi^{\prime}$ are the Eulerian perturbations in pressure, density and gravitational potential.
The solutions of this eigenvalue problem are given by a set of eigenvalues $\omega_{0,n,\ell,m}$ and eigenfunctions $\@vec{\delta r}_{0,n,\ell,m}$ that can be decomposed on the spherical harmonics basis as:
$\begin{split}\@vec{\delta r}_{0,n,\ell,m}&=\left[\xi_{0,r,\ell,n,m}(r)Y_{\ell}^{m}(\theta,\phi)\@vec{e}_{r}\right.\\ &\left.+\xi_{0,h,\ell,n,m}(r)\left(\frac{\partial Y_{\ell}^{m}}{\partial\theta}\@vec{e_{\theta}}+\frac{1}{\sin(\theta)}\frac{\partial Y_{\ell}^{m}}{\partial\varphi}\@vec{e_{\varphi}}\right)\right]\exp(-i\omega t)\,.\end{split}$ (3)
@@
$\begin{split}\@vec{\delta r}_{0,i}&=\left[\xi_{0,r,i}(r)Y_{\ell}^{m}(\theta,\phi)\@vec{e}_{r}\right.\\ &\left.+\xi_{0,h,i}(r)\left(\frac{\partial Y_{\ell}^{m}}{\partial\theta}\@vec{e_{\theta}}+\frac{1}{\sin(\theta)}\frac{\partial Y_{\ell}^{m}}{\partial\varphi}\@vec{e_{\varphi}}\right)\right]\exp(-i\omega t)\,.\end{split}$ (4)
@@
When including stellar rotation into the motion equation, the spherical-symmetry-induced degeneracy is lifted and the solutions to the eigenvalue problem depend on $m$. At leading order in $\Omega$, the angular rotation rate of the star, the eigenvalue problem becomes (e.g. 2010aste.book.....A; 2017A&A...605A..75D; 2025arXiv250926319A):
@@
$\omega^{2}\mathbb{I}\@vec{a}-\omega\mathbf{R}\@vec{a}-\mathbf{L}\@vec{a}=0\,$, (9)
where $\mathbb{I}$ is the identity matrix, $\mathbf{L}=\mathrm{diag}(\omega_{0,i}^{2})$, a\@vec{a} is the vector of the $a_{i}$ (i.e. the coefficients of the decomposition of $\@vec{\delta r}$ in the unperturbed base following Eq. (6)) and:
@@ 2.2 First order symmetric rotational effects: diagonal terms
may be asymptotically expressed using the seismic $\zeta$ function. For a given unperturbed mixed mode, the asymptotic expression of the $\zeta$-function provides the ratio between the mode inertia in the g-mode cavity and the total mode inertia (2013A&A...549A..75G, see also section 3.1.3 for the asymptotic expression of this quantity):
@@ 2.3 Near-degeneracy effects: non-diagonal terms
$\gamma_{c,ij}=\frac{1}{\sqrt{I_{0,i}I_{0,j}}}\int_{\mathrm{g}}K_{i,j}(r)\mathrm{d}r\,$, (18)
@@
$\gamma_{e,ij}=\frac{1}{\sqrt{I_{0,i}I_{0,j}}}\int_{\mathrm{p}}K_{i,j}(r)\mathrm{d}r\;.$ (19)
@@ 2.3.1 New asymptotic formula for the near degeneracy effects
Where we have used the trigonometric identity $\sin(a)\sin(b)=\frac{1}{2}\left(\cos(a-b)-\cos(a+b)\right)$. Using the change of variable111it is possible because in the g-mode cavity $\frac{N}{r}>0$ hence the integral is a strictly increasing function of $r$. This change of variables leads to $\mathrm{d}s=\frac{N}{r}\mathrm{d}r$
@@ 2.3.2 Validation of the new asymptotic formula for near degeneracy effects
This formulation is an approximation of (9) introduced by 2017A&A...605A..75D. Since they use this approximated form in their article, we choose to do the same when vetting our results in order to stay as close to the article as possible. This is the only part of the article in which we compute the splittings due to near-degeneracy effects with this approximation, in the rest of the article we solve the full quadratic eigenvalue problem (9).
@@ 3 Data analysis framework
To accomplish this, we employ a Monte-Carlo Markov Chain (MCMC) approach within a Bayesian framework to fit the oscillation modes in the PSD. The p- and mixed-mode patterns are created using the sloscillations code developed in 2019MNRAS.488..572K; 2023ApJ...954..152K, extended to incorporate the near-degeneracy effects based on Section 2. The synthetic PSD is generated following Sec. 3.1 below.
3.1 Synthetic PSD
3.1.1 Description of the set-up of the asymptotic model
The asymptotic model used in the fitting procedure predicts the full oscillation spectrum from a set of physically motivated parameters. For each angular degree $\ell\penalty 10000\ \in\penalty 10000\ \{0,1,2\}$, the model generates the underlying p- and g-mode frequencies from the asymptotic relations described below, which are then coupled through the mixed-mode formalism and corrected for rotational splitting at first order (for $\ell=1$) or through the full quadratic eigenvalue problem (for $\ell=2$).
The free parameters of the model therefore include the global p-mode pattern parameters ($\Delta\nu$, $\varepsilon$, $\delta\nu_{0,\ell}$, $\alpha$, $\beta_{0,\ell}$), the g-mode period-spacing and phase parameters ($\Delta\Pi_{\ell}$, $\varepsilon_{g,\ell}$), the p–g coupling strengths ($q_{\ell}$), and the core and envelope rotation rates ($\Omega_{\rm core}$, $\Omega_{\rm env}$), as well as the stellar inclination angle $\theta_{\rm inc}$.
Each predicted mode is assigned a Lorentzian profile with analytically prescribed linewidth and height, and these components are summed together with a synthetic granulation and white-noise background to produce a fully specified model of the PSD.
@@ 3.1.2 Pure pressure and gravity modes
$\nu_{{\rm g}_{n,\ell}}=\frac{1}{\Delta\Pi_{\ell}(n+\frac{1}{2}+\varepsilon_{\rm g})}\,$, (34)
@@ 3.1.3 Mixed modes with rotation
Unperturbed mixed mode frequencies are computed using equation (1979PASJ...31...87S):
@@
Then we introduce the effects of rotation on the unperturbed frequencies at first perturbative order for the $\ell=1$ components and solve the full rotating quadratic eigenvalue problem Eq. (9) with asymptotic matrix elements for $\ell=2$ (based on our asymptotic formulation Eq. (27)) modes using the slepc4py (10.1145/1089014.1089019; DALCIN20111124) Python library.
When computing the off diagonal elements of the rotation matrix, we use the asymptotic formula for $\zeta$ defined in Eq. 15 such that (2013A&A...549A..75G):
$\zeta(\nu)=\left[1+\frac{\nu^{2}\Delta\Pi_{\ell}}{\Delta\nu}\frac{1}{q\cos^{2}(\Theta_{p})+\frac{1}{q}\sin^{2}(\Theta_{p})}\right]^{-1}\,$, (37)
using the above definition of $\Theta_{p}$.
@@ 3.1.4 Modelling the linewidth, amplitudes and heights of the modes
This paragraph describes the equations and methodologies used to compute the linewidths amplitudes and heights of the modes in the sloscillations code by 2019MNRAS.488..572K for the sake of completeness and clarity. These parameters are fully derived from scaling relations and analytical expressions and are thus not fitted parameters in the model. We adopt effective-temperature-dependent scaling relations for the p mode linewidths, parameterised as a function of temperature, following the empirical prescriptions of corsaro_asteroseismology_2012; 2017ApJ...835..172L. Radial-mode amplitudes follow a Gaussian envelope centred on $\nu_{\rm max}$ with width tied to the envelope full width at half maximum and a peak amplitude that scales with global parameters ($\nu_{\rm max}$, $\Delta\nu$). The per-degree amplitudes are set by standard photometric visibilities as in 2019MNRAS.488..572K and, for non-radial multiplets, by geometric inclination ($\theta_{\mathrm{inc}}$) factors (2003ApJ...589.1009G). For mixed $\ell=1$ and $\ell=2$ modes, we modulate the nominal p mode amplitudes by $(1-\zeta)^{\frac{1}{2}}$ (2018A&A...618A.109M; Basu_Chaplin_2017). Mode heights in the power spectrum are computed from the time-domain amplitudes and the linewidths using the finite-observation-length correction of Fletcher_2006. For mixed modes, we use the mixed-mode generalisation that explicitly includes the inertia ratio $\zeta$ (Basu_Chaplin_2017), ensuring that g-dominated mixed modes appear with appropriately reduced height at fixed amplitude. Each mode contributes a Lorentzian profile to the PSD, centred at the asymptotic frequency and with the height and width described above.
The background of the synthetic PSD is modelled with the usual super-Lorentzian granulation components and white noise (e.g 2014A&A...570A..41K). It is generated in the time domain via a Gaussian-process (GP) representation of convection and granulation following 2019MNRAS.488..572K, where a kernel encapsulates the covariance of stochastic surface motions (e.g. 2017AJ....154..220F; 2019MNRAS.489.5764P; O’Sullivan_Aigrain_2024). This data is Fourier transformed, and the resulting PSD provides the granulation floor plus white noise that forms the background in the PSD. On top of this background, the oscillation spectrum is added, producing a self-consistent synthetic PSD where both the convective noise and p-mode power originate from the same underlying time-domain process. During the fitting procedure, the synthetic background is then normalised using the same procedure applied to the observed data (see Sec. 4.1).
3.2 Description of the Bayesian fitting procedure
@@ 3.2.1 Priors on p modes and $\ell=1$ mixed modes
3.2.2 Priors on the $\ell=2$ period spacing $\Delta\Pi_{2}$ and phase $\varepsilon_{g,2}$
@@
3.2.3 Priors on the $\ell=2$ modes coupling parameter $q_{2}$
We examine what constraints can be put on $q_{2}$, the coupling factor of $\ell=2$ modes, depending on the inclination angle of the star compared to the line of sight ($\theta_{\mathrm{inc}}$) (which we estimate by the preliminary fit of the $\ell=1$ modes – Appendix C). When the $\ell=2,\,m=0$ modes are visible, we use the technique described in Sec. 3.2.4. When the angle of inclination makes the $\ell=2$ and $m=0$ modes not visible, we use the result of Sec. 3.2.4.1.
3.2.4 Priors on the $\ell=2$ modes coupling parameter $q_{2}$ using the $m=0$ modes
@@
3.2.4.1 Priors on the $\ell=2$ modes coupling parameter $q_{2}$ using the $\ell=1$ mode coupling parameter $q_{1}$
@@
3.2.5 Likelihood adopted in the MCMC
In the MCMC, we adopt the likelihood appropriate for a power spectrum whose noise follows a $\chi^{2}$ distribution with two degrees of freedom (1984PhDT........34W; 1994A&A...289..649T; 2022A&A...663A.118B). The corresponding log-likelihood is
$\ln\mathcal{L}=-\sum_{i}\left[\frac{P_{i}}{M_{i}}+\ln M_{i}\right],$
where $P_{i}$ denotes the observed power and $M_{i}$ the model power at frequency bin $i$. The Bayesian inference procedure then adjusts the asymptotic parameters so as to maximise this likelihood, thereby ensuring that the resulting synthetic PSD matches the observed one as closely as possible. The different components of the model are presented in the paragraphs below.
3.2.6 Adopted fitting strategy
@@ 4.2 Results of the Bayesian asymptotic fitting
$\varepsilon$ $0.256\pm 0.004$ $0.425\pm 0.006$
$\delta\nu_{0,1}$ (µHz) $0.118\pm 0.02$ $0.346\pm 0.016$
$\delta\nu_{0,2}$ (µHz) $3.413\pm 0.021$ $2.182\pm 0.025$
@@
$\Delta\Pi_{1}$ (s) $111.180\pm 0.025$ $97.011\pm 0.008$
$q_{1}$ $0.379\pm 0.001$ $0.197\pm 0.001$
$\varepsilon_{g,1}$ $0.310\pm 0.006$ $0.279\pm 0.003$
$\Delta\Pi_{2}$ (s) $64.422\pm 0.007$ $56.015\pm 0.023$
$q_{2}$ $040\pm 0.001 0.0$$020\pm 0.001$
$\varepsilon_{g,2}$ $0.138\pm 0.004$ $0.135\pm 0.02$
${\Omega_{\mathrm{core}}}/{2\pi}$ (nHz) $781\pm 4$ $672\pm 8$
${\Omega_{\mathrm{env}}}/{2\pi}$ (nHz) $53\pm 6$ $49\pm 8$
$\theta_{\mathrm{inc}}$ (°) $85$ (fixed) $51\pm 3$